OpenStudy (anonymous):

Write the given equation in slope-intercept form, y = mx + b, by solving for y: 4x – 2y = 10 Answer y = 2x – 5 y = 2x + 5 y = -2x – 5 y = -2x + 5 The point (3, 0) is a solution of the equation: y = 2x − 11 Answer True False

OpenStudy (anonymous):

@aegidious

OpenStudy (anonymous):

u good at math?

OpenStudy (anonymous):

yeah, okay first you want to isolate the y value so to do that, you subtract the x value from both sides to get it onto the other side 4x-2y=10 -4x -4x -2y=10-4x

OpenStudy (anonymous):

make sense so far?

OpenStudy (anonymous):

ok yeah

OpenStudy (anonymous):

then you divide everything by -2 to get y all by itself

OpenStudy (anonymous):

ok i dont get it :/

OpenStudy (anonymous):

you need to get y=mx+b, right? well we have -2y=-4x+10

OpenStudy (anonymous):

we need to get that -2 off of the y, okay? that way y is all by itself

OpenStudy (anonymous):

oh ok now i get it

OpenStudy (anonymous):

so what answer did you get?

OpenStudy (anonymous):

is the answer (c) ?

OpenStudy (anonymous):

the 3rd one down? you don't have them lettered in the question so I don't know which is C

OpenStudy (anonymous):

is this the answer y = -2x – 5

OpenStudy (anonymous):

the third one

OpenStudy (anonymous):

you forget that when you're dividing negatives, they become positive. -4x is a negative divided by a negative -2

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

so if they become positive numbers then i think the answer is y=2x+5

OpenStudy (anonymous):

mmmm the negative -4x becomes a positive 2x, yes BUT a positive divided by a negative becomes negative. The positive 10 divided by a negative then becomes.......what? ;)

OpenStudy (anonymous):

u know it would be very nice if u would just give me the answer:]

OpenStudy (anonymous):

ok the answer has to be y=2x-5 right....

OpenStudy (anonymous):

YES! :)

OpenStudy (anonymous):

lol ty so u gonna help me with the second Q?

OpenStudy (anonymous):

and then for the second one, all you do is plug 3 into the x value. If you calculate your y as being 0, then the answer is true. If you get any other number for y, then it's false

OpenStudy (anonymous):

the answer is -5 so is it false?

OpenStudy (anonymous):

Exactly :)

OpenStudy (anonymous):

yay -pats myself on the back- :]

OpenStudy (anonymous):

Calculate the slope of the points (-2, -1) and (3, 4). Answer 1 2 3 4 wb this Q

OpenStudy (anonymous):

give a medal if you like my help ;)

OpenStudy (anonymous):

ok but if u just leave after i give it to u im gonna end ur life :]

OpenStudy (anonymous):

come at me bro!!! XD

OpenStudy (anonymous):

lolol okay, do you know rise over run?

OpenStudy (anonymous):

oh trust me i will

OpenStudy (anonymous):

did just gimmi the answer and i'll fan u

OpenStudy (anonymous):

the slope is 1

OpenStudy (anonymous):

good boy :]

OpenStudy (anonymous):

Solve the following system of linear equations. 3x + 2y = 10 2x + 3y = 15/2 Answer No solution y = (-3/2)x + 5 x = 3, y = -1/2 x = 3, y = 1/2<----this is my answer

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

nuh huh, can only do so much ;) If you want my help, open another question XD

OpenStudy (anonymous):

i hate u ok give me a second